OpenStudy (countrygirl1431):

Please help!!! Type in expanded form. Use parentheses to indicate multiplication. (2r ) 5

jimthompson5910 (jim_thompson5910):

The question is \[\Large (2r)^5\] or no?

OpenStudy (countrygirl1431):

yeah

jimthompson5910 (jim_thompson5910):

raise each factor inside the parenthesis to the 5th power to get \[\Large (2r)^5 = 2^5*r^5 = ??\]

OpenStudy (countrygirl1431):

i still dont get it lol

jimthompson5910 (jim_thompson5910):

what is 2^5 equal to?

jimthompson5910 (jim_thompson5910):

2^5 = 2*2*2*2*2 = ?? the exponent 5 means there are 5 copies of "2" being multiplied

OpenStudy (countrygirl1431):

right i understand that much

jimthompson5910 (jim_thompson5910):

you should find that 2^5 = 32, agreed?

OpenStudy (countrygirl1431):

yeah

jimthompson5910 (jim_thompson5910):

So, \[\Large (2r)^5 = 2^5*r^5 = 32r^5\]

OpenStudy (countrygirl1431):

oh ok i see i got it now thanks

jimthompson5910 (jim_thompson5910):

no problem